# Brewing Vodka: Distilling Pure Knowledge for Lightweight Threat Detection in Audit Logs

### v1

@@ Abstract.
To address these challenges, we propose Vodka, a lightweight threat detection system built on a knowledge distillation framework, capable of node-level detection within audit log provenance graphs. Specifically, Vodka applies graph Laplacian regularization to reduce neighbor noise, obtaining smoothed and denoised graph signals. Subsequently, Vodka employs a teacher model based on GNNs to extract knowledge, which is then distilled into a lightweight student model. The student model is designed as a trainable combination of a feature transformation module and a personalized PageRank random walk label propagation module, with the former capturing feature knowledge and the latter learning label and structural knowledge. After distillation, the student model benefits from the knowledge of the teacher model to perform precise threat detection. Finally, Vodka reconstructs attack paths from anomalous nodes, providing insight into the attackers’ strategies. We evaluate Vodka through extensive experiments on three public datasets and compare its performance against several state-of-the-art IDS solutions. The results demonstrate that Vodka achieves outstanding detection accuracy across all scenarios and the detection time is 1.4 to 5.2 times faster than the current state-of-the-art methods.
@@ 1. Introduction
Thus, we propose Vodka, a lightweight APTs detection method that achieves both high detection accuracy and computational efficiency. Specifically, Vodka incorporates the following key functionalities: (1) Vodka constructs a provenance graph from audit logs that include various APTs attacks and uses Word2Vec to assign initial features to nodes. To address the issue of neighbor noise, Vodka designs a graph Laplacian regularization-based graph signal denoising method tailored to the provenance graph. This algorithm smooths node signals without disrupting the original graph topology. (2) Vodka introduces a knowledge distillation framework where, during training, knowledge from a pre-trained large GNN teacher model is distilled into a lightweight student model. The lightweight student model is then used for anomalous node detection, significantly reducing detection costs. (3) For the student model, Vodka proposes a new hybrid mechanism that combines the Feature Transformation (FT) mechanism with a personalized PageRank random walk label propagation (PRL). This allows more effective learning of prior knowledge from features, labels, and structures while utilizing the GNN knowledge from the teacher model. (4) After detecting anomalous nodes, Vodka uses a community division algorithm to identify malicious communities.
We performed a comprehensive evaluation of Vodka using datasets widely adopted by the research community, including StreamSpot(The, 2016), Unicorn Wget(Han et al., 2020), and DARPA-E3(DAR, 2020). We first analyzed its detection performance and compared it against various existing baselines, demonstrating that Vodka consistently outperforms almost all these systems under the same evaluation metrics. Furthermore, we analyzed Vodka’s detection costs and explored its potential as a real-time detection system. In addition, hyperparameter analysis and ablation experiments on Vodka highlight the irreplaceability of its key parameters and main components. Lastly, we verified Vodka’s inherent robustness against adversarial attacks. The primary contributions of our work are as follows:
@@
  • We propose Vodka, a general APTs detection method based on knowledge distillation, capable of transferring GNN knowledge from a large teacher model to a lightweight student model, thereby reducing resource and time costs during detection.
@@
  • We comprehensively evaluate Vodka ’s performance by implementing it on three widely used datasets and conducting various experiments to verify its effectiveness. Experimental results demonstrate that Vodka outperforms most existing methods while incurring lower detection costs.
@@ 2.1. Motivating Example
Vodka (ours) Node
@@ 2.2. Limitations to Existing Solutions
Learning-based detection methods are currently the most promising in the field of threat detection. Their core principle is to build anomaly models by learning from large datasets and then infer whether anomalous behavior represents a malicious attack. Some of the most notable recent works include: MAGIC(Jia et al., 2024), which performs efficient detection by masking the graph and reducing costs; KAIROS(Cheng et al., 2023), which identifies four dimensions of intrusion detection systems and uses an encoder-decoder architecture to detect attacks and reconstruct attack chains; THREATRACE(Wang et al., 2022), the first detector to propose behavior pattern construction for each type of node in the provenance graph to achieve node-level detection; ShadeWatcherShadeWatcher(Zengy et al., 2022), which introduces recommendation system techniques into provenance graph detection and achieves edge-level detection; Unicorn(Han et al., 2020), which builds a provenance graph enriched with semantics and historical information to capture long-term stealthy attacks. Although these works demonstrate excellent performance in threat detection, they still face limitations when tested in real-world environments. These limitations are what Vodka aims to address, as shown in Table 1.
@@ 2.3. Threat Model
Although the attackers’ behavior is highly covert, audit logs capture complete, traceable evidence and footprints of their activities. As in most of the existing works in this field(Wang et al., 2022; Han et al., 2020; Rehman et al., 2024), we assume that the underlying operating system, audit framework, and system hardware are intact and trustworthy(Bates et al., 2015; Pasquier et al., 2017). Additionally, we assume that attackers cannot directly tamper with the contents of the audit logs, ensuring that the provenance graph constructed by Vodka remains reliable(Paccagnella et al., 2020).
@@ 3. Methodology
Figure 2. The framework of Vodka.
In this section, we provide a detailed overview of Vodka’s overall design framework and the design of each module. The framework consists of the following components: ①Graph Construction. ②Log Distillation. ③Threat Detection. ④Attack Reconstruction. The overall framework is illustrated in Figure 2.
@@ 3.1.1. Provenance Graph Construction
Vodka processes host audit logs from various sources(such as Windows ETW, Linux Audit and CamFlow(Pasquier et al., 2017)) and transforms them into a graph structure known as a Provenance Graph (PG). In this graph, nodes represent system entities, and edges denote the interactions between these entities.
Since system logs contain rich attributes related to various system entities, Vodka encodes these attributes into a vector space to serve as initial input for model learning. We use Word2Vec(Ma and Zhang, 2015), a neural network-based technique effective at learning word vector representations, which generates dense, low-dimensional vectors for each word. In our cataloging process, we use attributes such as file paths for files, process names for process nodes, system entity types, and network IP addresses for sockets to generate feature vectors. These are combined with semantic attributes and system call types within one hop of the neighbors to form sentences. During training, Vodka captures semantic relationships between words within a fixed length, outputting embedded vectors as the raw signal for subsequent modules.
@@ 3.1.2. Neighbor Denoising
Given that in the provenance graph constructed from audit logs, malicious nodes (i.e., system entities manipulated by attackers) are rare, and adversarial activities—such as deceptive access to benign nodes or the manipulation of compromised benign nodes to perform misleading actions—are common, there exists a significant issue of neighbor noise for anomaly detection tasks. To address this, Vodka incorporates a graph Laplacian regularization-based signal denoising method, which is designed to leverage the graph’s topological structure to smooth signals and reduce the impact of noise.
Specifically, Vodka defines the weights based on the similarity and connection relationships between entity nodes and combines them into a weight matrix $W$. By calculating the node degree matrix $D$, we can then derive the Laplacian matrix $L=D-W$. This enables the formulation of a convex quadratic optimization problem to minimize noise and smooth signals as follows:
@@
After the neighbor denoising process, we obtain denoised and smoothed signal vectors. It is worth noting that Vodka’s neighbor denoising algorithm does not alter the graph’s topology in any way, ensuring that no new noise is introduced due to changes in the structure. This preserves the integrity of the original graph topology while effectively reducing the noise, allowing for more accurate and reliable detection in subsequent steps.
@@ 3.2. Log Distillation
In this section, we describe the knowledge distillation mechanism designed for Vodka ’s log provenance graph, which explains the training process of the teacher and student models and how knowledge is distilled from the large-scale teacher model to the lightweight student model.
@@ 3.2.1. Teacher Model
The teacher model is the cornerstone of Vodka ’s APTs threat detection system, essential for reducing detection time costs. The teacher model is pre-trained before real-time detection occurs. For Vodka ’s teacher model, we employ various large-scale GNN models, such as GCN(Kipf and Welling, 2016), GAT(Veličković et al., 2017), GraphSAGE(Hamilton et al., 2017), and SGC(Wu et al., 2019). These GNN models are well-suited for leveraging the complex interactions between system entities found in audit logs and capturing causal connections and semantic information across entities. This characteristic is commonly utilized by the most advanced IDS today.
Unlike other IDS, Vodka ’s teacher model is trained using supervised labels on system entities fed into the GNN model, producing both the predicted result $f_{t}hr$ and soft labels. Soft labels represent the aggregated anomaly probabilities for all test nodes and serve as an important self-supervised signal for the subsequent student model. By pre-training the teacher model on the large-scale graph topology derived from the log data and generating soft labels, Vodka ensures robust support for real-time detection.
@@ 3.2.2. Student Model
The core requirement for Vodka ’s student model is to compress detection time while maintaining high accuracy and effectively utilizing prior knowledge. Thus, the student model is designed to be lightweight and efficient, built on two foundational mechanisms: 1) feature transformation for individual nodes, and 2) label propagation(Iscen et al., 2019) between nodes. The former focuses on enhancing the model’s ability to fit based on each node’s intrinsic features, while the latter propagates label information across nodes.
Vodka ’s student model integrates both mechanisms, combining Feature Transformation (FT) and Personalized PageRank Random walk Label propagation (PRL). Feature Transformation enables each node to adjust and refine its feature representation to improve detection accuracy, while PRL helps spread labels effectively across the graph to capture anomaly information in a more efficient and informed manner. This hybrid approach allows the student model to leverage the knowledge distilled from the teacher model while significantly reducing detection time and computational costs.
1. Feature Transformation We use a simple Multilayer Perceptron (MLP)(Tang et al., 2015) for feature transformation to perform nonlinear transformations on the input node features, enhancing the model’s ability to fit complex data. In general, MLP takes the raw node features as input and outputs the predicted probability for malicious nodes. Vodka employs a two-layer MLP.
@@ 3.2.3. Optimization Objective
Figure 3. Vodka working modes.
@@ 3.3. Threat Detection
Our student model operates in two modes: a training mode and a detection mode. As shown in Figure 3, during the knowledge distillation process, we transfer knowledge from the complex teacher model to the lightweight student model. In this phase, the student model is in training mode, focusing on learning the distilled knowledge from the teacher model. Once the knowledge distillation step is complete, Vodka employs the trained student model as a detection model, which infers and classifies anomalous nodes from the provenance graph constructed from logs. At this point, the student model switches to detection mode. The detection mode workflow is as follows: we input the provenance graph (with initial features) to the student model. After lightweight $K$-layer PRL and feature transformation steps, the student model predicts an anomaly score for each node $v$. The student model efficiently passes messages and aggregates information across the input graph, classifying nodes based on the learned lightweight feature representations. The anomaly score output for each node is then compared to a pre-set threshold, and nodes exceeding this threshold are flagged as potential malicious nodes and collected into the malicious node set.
@@ 3.4. Attack Reconstruction
After threat detection, security personnel may have a set of detected malicious nodes provided by Vodka, but these nodes are often dispersed across the graph, making it difficult to directly derive the attacker’s path. To alleviate the burden on security analysts, Vodka reconstructs a more detailed attack trace, capturing the complete attack chain without relying on prior knowledge. Specifically, Vodka leverages Infomap(Edler et al., 2017), an algorithm based on random walks that simulates how information propagates between nodes and divides the graph into tightly connected communities. Malicious nodes and their one-hop neighbors are assigned higher weights to distinguish them from benign node connections. Next, information flow propagation is performed, where information spreads across the graph according to weighted probabilities, merging nodes to form communities.
@@ 4. Evaluation
In this section, we will conduct various experiments to validate the advantages of Vodka and answer the following research questions:
  • RQ1: How does Vodka ’s detection performance compare to state-of-the-art methods?
  • RQ2: How effective are the main components of Vodka?
  • RQ3: How do Vodka ’s hyperparameters affect its detection performance?
  • RQ4: What are the cost overheads of Vodka as a detection system?
  • RQ5: How robust is Vodka against adversarial attacks?
@@ 4.1.1. Datasets
DARPA-E3 Dataset(DAR, 2020): This dataset is part of the DARPA TC program’s third evaluation, containing several sub-datasets. We selected the Cadets, Theia, and Trace datasets to evaluate Vodka, using the same ground truth labels as ThreaTrace. Detailed descriptions of these datasets can be found in the appendix.
@@ 4.1.3. Baselines
To perform a comprehensive performance evaluation, we compared Vodka with the best detection methods. Unfortunately, some methods were excluded due to challenges in reproducibility or incompatibility. For example, KAIROS(Cheng et al., 2023) could not be compared as it uses a time window-based approach that differs from other methods. We selected the following detection methods as our baselines:
@@ 4.1.4. Implementation
Details We implemented Vodka ’s prototype in approximately 2800 lines of Python3.11 code. For log processing, we used a log parser to convert audit log files from various sources into JSON format and then preprocess them. The development environment was PyTorch(Paszke et al., 2019), and the model implementation was supported by the Deep Graph Library (DGL)(Wang, 2019). Detailed teacher and student model parameters are provided in the A Appendix.

### v2

# Winemaking: Extracting Essential Insights for Efficient Threat Detection in Audit Logs

@@ Abstract.
To address these challenges, we propose Winemaking, a lightweight threat detection system built on a knowledge distillation framework, capable of node-level detection within audit log provenance graphs. Specifically, Winemaking applies graph Laplacian regularization to reduce neighbor noise, obtaining smoothed and denoised graph signals. Subsequently, Winemaking employs a teacher model based on GNNs to extract knowledge, which is then distilled into a lightweight student model. The student model is designed as a trainable combination of a feature transformation module and a personalized PageRank random walk label propagation module, with the former capturing feature knowledge and the latter learning label and structural knowledge. After distillation, the student model benefits from the knowledge of the teacher model to perform precise threat detection. Finally, Winemaking reconstructs attack paths from anomalous nodes, providing insight into the attackers’ strategies. We evaluate Winemaking through extensive experiments on three public datasets and compare its performance against several state-of-the-art IDS solutions. The results demonstrate that Winemaking achieves outstanding detection accuracy across all scenarios and the detection time is 1.4 to 5.2 times faster than the current state-of-the-art methods.
@@ 1. Introduction
Thus, we propose Winemaking, a lightweight APTs detection method that achieves both high detection accuracy and computational efficiency. Specifically, Winemaking incorporates the following key functionalities: (1) Winemaking constructs a provenance graph from audit logs that include various APTs attacks and uses Word2Vec to assign initial features to nodes. To address the issue of neighbor noise, Winemaking designs a graph Laplacian regularization-based graph signal denoising method tailored to the provenance graph. This algorithm smooths node signals without disrupting the original graph topology. (2) Winemaking introduces a knowledge distillation framework where, during training, knowledge from a pre-trained large GNN teacher model is distilled into a lightweight student model. The lightweight student model is then used for anomalous node detection, significantly reducing detection costs. (3) For the student model, Winemaking proposes a new hybrid mechanism that combines the Feature Transformation (FT) mechanism with a personalized PageRank random walk label propagation (PRL). This allows more effective learning of prior knowledge from features, labels, and structures while utilizing the GNN knowledge from the teacher model. (4) After detecting anomalous nodes, Winemaking uses a community division algorithm to identify malicious communities.
We performed a comprehensive evaluation of Winemaking using datasets widely adopted by the research community, including StreamSpot(The, 2016), Unicorn Wget(Han et al., 2020), and DARPA-E3(DAR, 2020). We first analyzed its detection performance and compared it against various existing baselines, demonstrating that Winemaking consistently outperforms almost all these systems under the same evaluation metrics. Furthermore, we analyzed Winemaking’s detection costs and explored its potential as a real-time detection system. In addition, hyperparameter analysis and ablation experiments on Winemaking highlight the irreplaceability of its key parameters and main components. Lastly, we verified Winemaking’s inherent robustness against adversarial attacks. The primary contributions of our work are as follows:
@@
  • We propose Winemaking, a general APTs detection method based on knowledge distillation, capable of transferring GNN knowledge from a large teacher model to a lightweight student model, thereby reducing resource and time costs during detection.
@@
  • We comprehensively evaluate Winemaking ’s performance by implementing it on three widely used datasets and conducting various experiments to verify its effectiveness. Experimental results demonstrate that Winemaking outperforms most existing methods while incurring lower detection costs.
@@ 2.1. Motivating Example
Winemaking (ours) Node
@@ 2.2. Limitations to Existing Solutions
Learning-based detection methods are currently the most promising in the field of threat detection. Their core principle is to build anomaly models by learning from large datasets and then infer whether anomalous behavior represents a malicious attack. Some of the most notable recent works include: MAGIC(Jia et al., 2024), which performs efficient detection by masking the graph and reducing costs; KAIROS(Cheng et al., 2023), which identifies four dimensions of intrusion detection systems and uses an encoder-decoder architecture to detect attacks and reconstruct attack chains; THREATRACE(Wang et al., 2022), the first detector to propose behavior pattern construction for each type of node in the provenance graph to achieve node-level detection; ShadeWatcherShadeWatcher(Zengy et al., 2022), which introduces recommendation system techniques into provenance graph detection and achieves edge-level detection; Unicorn(Han et al., 2020), which builds a provenance graph enriched with semantics and historical information to capture long-term stealthy attacks. Although these works demonstrate excellent performance in threat detection, they still face limitations when tested in real-world environments. These limitations are what Winemaking aims to address, as shown in Table 1.
@@ 2.3. Threat Model
Although the attackers’ behavior is highly covert, audit logs capture complete, traceable evidence and footprints of their activities. As in most of the existing works in this field(Wang et al., 2022; Han et al., 2020; Rehman et al., 2024), we assume that the underlying operating system, audit framework, and system hardware are intact and trustworthy(Bates et al., 2015; Pasquier et al., 2017). Additionally, we assume that attackers cannot directly tamper with the contents of the audit logs, ensuring that the provenance graph constructed by Winemaking remains reliable(Paccagnella et al., 2020).
@@ 3. Methodology
Figure 2. The framework of Winemaking.
In this section, we provide a detailed overview of Winemaking’s overall design framework and the design of each module. The framework consists of the following components: ①Graph Construction. ②Log Distillation. ③Threat Detection. ④Attack Reconstruction. The overall framework is illustrated in Figure 2.
@@ 3.1.1. Provenance Graph Construction
Winemaking processes host audit logs from various sources(such as Windows ETW, Linux Audit and CamFlow(Pasquier et al., 2017)) and transforms them into a graph structure known as a Provenance Graph (PG). In this graph, nodes represent system entities, and edges denote the interactions between these entities.
Since system logs contain rich attributes related to various system entities, Winemaking encodes these attributes into a vector space to serve as initial input for model learning. We use Word2Vec(Ma and Zhang, 2015), a neural network-based technique effective at learning word vector representations, which generates dense, low-dimensional vectors for each word. In our cataloging process, we use attributes such as file paths for files, process names for process nodes, system entity types, and network IP addresses for sockets to generate feature vectors. These are combined with semantic attributes and system call types within one hop of the neighbors to form sentences. During training, Winemaking captures semantic relationships between words within a fixed length, outputting embedded vectors as the raw signal for subsequent modules.
@@ 3.1.2. Neighbor Denoising
Given that in the provenance graph constructed from audit logs, malicious nodes (i.e., system entities manipulated by attackers) are rare, and adversarial activities—such as deceptive access to benign nodes or the manipulation of compromised benign nodes to perform misleading actions—are common, there exists a significant issue of neighbor noise for anomaly detection tasks. To address this, Winemaking incorporates a graph Laplacian regularization-based signal denoising method, which is designed to leverage the graph’s topological structure to smooth signals and reduce the impact of noise.
Specifically, Winemaking defines the weights based on the similarity and connection relationships between entity nodes and combines them into a weight matrix $W$. By calculating the node degree matrix $D$, we can then derive the Laplacian matrix $L=D-W$. This enables the formulation of a convex quadratic optimization problem to minimize noise and smooth signals as follows:
@@
After the neighbor denoising process, we obtain denoised and smoothed signal vectors. It is worth noting that Winemaking’s neighbor denoising algorithm does not alter the graph’s topology in any way, ensuring that no new noise is introduced due to changes in the structure. This preserves the integrity of the original graph topology while effectively reducing the noise, allowing for more accurate and reliable detection in subsequent steps.
@@ 3.2. Log Distillation
In this section, we describe the knowledge distillation mechanism designed for Winemaking ’s log provenance graph, which explains the training process of the teacher and student models and how knowledge is distilled from the large-scale teacher model to the lightweight student model.
@@ 3.2.1. Teacher Model
The teacher model is the cornerstone of Winemaking ’s APTs threat detection system, essential for reducing detection time costs. The teacher model is pre-trained before real-time detection occurs. For Winemaking ’s teacher model, we employ various large-scale GNN models, such as GCN(Kipf and Welling, 2016), GAT(Veličković et al., 2017), GraphSAGE(Hamilton et al., 2017), and SGC(Wu et al., 2019). These GNN models are well-suited for leveraging the complex interactions between system entities found in audit logs and capturing causal connections and semantic information across entities. This characteristic is commonly utilized by the most advanced IDS today.
Unlike other IDS, Winemaking ’s teacher model is trained using supervised labels on system entities fed into the GNN model, producing both the predicted result $f_{t}hr$ and soft labels. Soft labels represent the aggregated anomaly probabilities for all test nodes and serve as an important self-supervised signal for the subsequent student model. By pre-training the teacher model on the large-scale graph topology derived from the log data and generating soft labels, Winemaking ensures robust support for real-time detection.
@@ 3.2.2. Student Model
The core requirement for Winemaking ’s student model is to compress detection time while maintaining high accuracy and effectively utilizing prior knowledge. Thus, the student model is designed to be lightweight and efficient, built on two foundational mechanisms: 1) feature transformation for individual nodes, and 2) label propagation(Iscen et al., 2019) between nodes. The former focuses on enhancing the model’s ability to fit based on each node’s intrinsic features, while the latter propagates label information across nodes.
Winemaking ’s student model integrates both mechanisms, combining Feature Transformation (FT) and Personalized PageRank Random walk Label propagation (PRL). Feature Transformation enables each node to adjust and refine its feature representation to improve detection accuracy, while PRL helps spread labels effectively across the graph to capture anomaly information in a more efficient and informed manner. This hybrid approach allows the student model to leverage the knowledge distilled from the teacher model while significantly reducing detection time and computational costs.
1. Feature Transformation We use a simple Multilayer Perceptron (MLP)(Tang et al., 2015) for feature transformation to perform nonlinear transformations on the input node features, enhancing the model’s ability to fit complex data. In general, MLP takes the raw node features as input and outputs the predicted probability for malicious nodes. Winemaking employs a two-layer MLP.
@@ 3.2.3. Optimization Objective
Figure 3. Winemaking working modes.
@@ 3.3. Threat Detection
Our student model operates in two modes: a training mode and a detection mode. As shown in Figure 3, during the knowledge distillation process, we transfer knowledge from the complex teacher model to the lightweight student model. In this phase, the student model is in training mode, focusing on learning the distilled knowledge from the teacher model. Once the knowledge distillation step is complete, Winemaking employs the trained student model as a detection model, which infers and classifies anomalous nodes from the provenance graph constructed from logs. At this point, the student model switches to detection mode. The detection mode workflow is as follows: we input the provenance graph (with initial features) to the student model. After lightweight $K$-layer PRL and feature transformation steps, the student model predicts an anomaly score for each node $v$. The student model efficiently passes messages and aggregates information across the input graph, classifying nodes based on the learned lightweight feature representations. The anomaly score output for each node is then compared to a pre-set threshold, and nodes exceeding this threshold are flagged as potential malicious nodes and collected into the malicious node set.
@@ 3.4. Attack Reconstruction
After threat detection, security personnel may have a set of detected malicious nodes provided by Winemaking, but these nodes are often dispersed across the graph, making it difficult to directly derive the attacker’s path. To alleviate the burden on security analysts, Winemaking reconstructs a more detailed attack trace, capturing the complete attack chain without relying on prior knowledge. Specifically, Winemaking leverages Infomap(Edler et al., 2017), an algorithm based on random walks that simulates how information propagates between nodes and divides the graph into tightly connected communities. Malicious nodes and their one-hop neighbors are assigned higher weights to distinguish them from benign node connections. Next, information flow propagation is performed, where information spreads across the graph according to weighted probabilities, merging nodes to form communities.
@@ 4. Evaluation
In this section, we will conduct various experiments to validate the advantages of Winemaking and answer the following research questions:
  • RQ1: How does Winemaking ’s detection performance compare to state-of-the-art methods?
  • RQ2: How effective are the main components of Winemaking?
  • RQ3: How do Winemaking ’s hyperparameters affect its detection performance?
  • RQ4: What are the cost overheads of Winemaking as a detection system?
  • RQ5: How robust is Winemaking against adversarial attacks?
@@ 4.1.1. Datasets
DARPA-E3 Dataset(DAR, 2020): This dataset is part of the DARPA TC program’s third evaluation, containing several sub-datasets. We selected the Cadets, Theia, and Trace datasets to evaluate Winemaking, using the same ground truth labels as ThreaTrace. Detailed descriptions of these datasets can be found in the appendix.
@@ 4.1.3. Baselines
To perform a comprehensive performance evaluation, we compared Winemaking with the best detection methods. Unfortunately, some methods were excluded due to challenges in reproducibility or incompatibility. For example, KAIROS(Cheng et al., 2023) could not be compared as it uses a time window-based approach that differs from other methods. We selected the following detection methods as our baselines:
@@ 4.1.4. Implementation
Details We implemented Winemaking ’s prototype in approximately 2800 lines of Python3.11 code. For log processing, we used a log parser to convert audit log files from various sources into JSON format and then preprocess them. The development environment was PyTorch(Paszke et al., 2019), and the model implementation was supported by the Deep Graph Library (DGL)(Wang, 2019). Detailed teacher and student model parameters are provided in the A Appendix.